The half-life of radon-222 is 3.8 days. How much of a 100.o g sample is left after 15.2 days


1
Expert's answer
2021-12-07T04:33:04-0500

If 4 half-lives have elapsed, 1/2×1/2×1/2×1/2 = 1/16 of the original quantity of 222Rn will remain.

Thus 100 g × 1/16 = 6.25 g